# Learning to Recover: Dynamic Reward Shaping with Wheel-Leg Coordination for Fallen Robots

### v1

@@ IV-D Cross-Platform Validation on Different Robots
To validate the cross-platform generalization capability of the DS strategy, we deployed it on the Unitree Go2-W[11] platform with significantly different configuration parameters and compared its performance with that of the baseline method. The experimental results indicate that despite hardware differences such as mass distribution and joint drive limits, the DS strategy maintained robust recovery performance and demonstrated the advantages of the wheel–leg collaboration mechanism. As illustrated in Fig11, with the wheels actuated, the Go2-W robot utilized active wheel rotation to buffer its posture, thereby significantly reducing the magnitude of torque adjustments required at the hip joints. Its recovery success rate improved by $3.7$ percentage points compared to the baseline strategy, see TableII, and the average joint torque was reduced by $26.2\%$ relative to the fixed-wheel configuration (see TableIII and Fig.8).
These findings are consistent with the experimental results on the KYON platform, indicating that the DS strategy consistently generates optimized action sequences tailored to local degrees of freedom via dynamic reward shaping, and that the wheeled assistance contact force regulation mechanism effectively reduces joint loads.
@@
We also performed a quantitative analysis of the joint motion parameters for both robots. As shown in Fig9, by calculating the mean and standard deviation of the peak velocities for each joint, we verified that the driving commands generated by our control strategy adhered to the maximum velocity thresholds specified in the robot description files (KYON: $7.6rad/s$ and Unitree Go2-W: $20.3rad/s$). Furthermore, we also evaluated the mean and standard deviation of the torques for each joint, as illustrated in Fig.10. For the Unitree Go2-W robot, most joints exhibit a maximum torque of $23.7N\cdot m$, with the calf joint(KFE) reaching up to $35.55N\cdot m$. In contrast, all joints of the KYON robot achieves a maximum joint torque of $185N\cdot m$. All movements stay within the allowable joint range limits, confirming that the actions generated by our simulation-based control policy are indeed feasible for real-world experiments.
@@
Platform Configuration Avg Torque (N·m) Reduction (%)
@@
IV-E Recovery Adaptability on Unseen Terrains
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Notably, the policy exhibits partial zero-shot generalization: the robot maintains stability through controlled wheel-leg coordination, despite the absence of explicit terrain sensing. We compared the recovery success rates of the DS strategy and the baseline strategy deployed on the KYON robot in non-flat environments, defining recovery as having a joint angle deviation from the default configuration of no more than $0.5rad$ and a base orientation error—calculated as the sum of differences between the robot’s projected gravity vector and the ideal direction $\left[0,0,-1\right]$ —of less than $0.1$, with the success rates shown in the tableIV.
However, recovery success rates decline significantly under these conditions, with frequent secondary falls observed—particularly on stair terrains. These undesirable behaviors, which undermine real-world deployment, can be attributed to two unresolved challenges: first, the policy’s reactive adjustments lack awareness of terrain geometry; and second, it fails to optimally adapt its recovery posture to local terrain features.
@@
Figure 12: Evaluation of the DS-policy on the KYON Robot in irregular terrain.
@@ V CONCLUSION
This paper presents a Episode-based Dynamic Reward Shaping for wheel-legged robots to achieve robust and adaptive recovery from fallen states. By integrating dynamic reward shaping and curriculum learning, the proposed method effectively addresses the exploration-convergence dilemma inherent in sparse-reward reinforcement learning. Experimental results demonstrate that DS achieves a recovery success rate of over $97\%$ across diverse platforms, outperforming baseline methods by 3 percentage points. The key innovation lies in the synergy between wheel-assisted motion and leg articulation: wheels provide rapid centroidal adjustment through controlled rolling, while legs enable precise posture stabilization, collectively reducing joint torque demands by $15-26\%$ compared to fixed-wheel configurations. Cross-platform validation on robots with distinct kinematic and dynamic parameters (KYON and Unitree Go2-W[11]) further confirms the generalization capability of the framework, where DS autonomously adapts to hardware-specific constraints without manual tuning. These findings highlight the potential of wheel-leg coordination as a universal strategy for recovery control in hybrid locomotion systems.
Future work will focus on extending the framework to handle more complex terrains (e.g., slopes, gravel) and integrating real-world sensor noise models to bridge the sim-to-real gap. Additionally, the principles of dynamic reward shaping could be applied to other discontinuous-contact tasks, such as multi-robot collaborative recovery or manipulation-oriented posture adjustment.
@@ ACKNOWLEDGMENT
We would like to thank Andrea Patrizi, Despoina Maligianni, Rui Dai, Yifang Zhang Maolin Lei, Jingcheng Jiang, Kuanqi Cai, and Carlo Rizzardo for their useful discussions.

### v2

@@ IV-D Cross-Platform Validation on Different Robots
To validate the cross-platform generalization capability of the ED strategy, we deployed it on the Unitree Go2-W[8] platform with significantly different configuration parameters and compared its performance to a baseline in which the ED modules were ablated to isolate the effect of our proposed strategy. The experimental results indicate that despite hardware differences such as mass distribution and joint drive limits, the ED strategy maintained robust recovery performance and demonstrated the advantages of the wheel–leg collaboration mechanism. As illustrated in Fig. 11, with the wheels actuated, the Go2-W robot utilized active wheel rotation to buffer its posture, thereby significantly reducing the magnitude of torque adjustments required at the hip joints. Its recovery success rate improved by $3.7$ percentage points compared to the baseline strategy, see Table II, and the average joint torque was reduced by $26.2\%$ relative to the fixed-wheel configuration (see Table III and Fig. 8).
These findings are consistent with the experimental results on the KYON platform, indicating that the ED strategy consistently generates optimized action sequences tailored to local degrees of freedom via dynamic reward shaping, and that the wheeled assistance contact force regulation mechanism effectively reduces joint loads.
@@
We also performed a quantitative analysis of the joint motion parameters for both robots. As shown in Fig. 9, by calculating the mean and standard deviation of the peak velocities for each joint, we verified that the driving commands generated by our control strategy adhered to the maximum velocity thresholds specified in the robot description files (KYON: $7.6rad/s$ and Unitree Go2-W: $20.3rad/s$). Furthermore, we also evaluated the mean and standard deviation of the torques for each joint, as illustrated in Fig. 10. For the Unitree Go2-W robot, most joints exhibit a maximum torque of $23.7N\cdot m$, with the calf joint(KFE) reaching up to $35.55N\cdot m$. In contrast, all joints of the KYON robot achieves a maximum joint torque of $185N\cdot m$. All movements stay within the allowable joint range limits, confirming that the actions generated by our simulation-based control policy are indeed feasible for real-world experiments.
@@
Platform Configuration Avg Torque ($N\cdot m$) Reduction (%)
@@
IV-E Recovery under Repeated Perturbations
To accelerate training, we adopted a fixed-step recovery setting in simulation. In real-world tests, once the robot reached a standing posture, we applied multiple manual perturbations without reset. The policy autonomously recovered after each perturbation, Fig. 12 (a)–(g); when the initial recovery was insufficient, it immediately corrected the motion and performed a second recovery, Fig. 12 (h)–(j). These results indicate that, at test time, recovery is sustained and not constrained by the fixed step used during training.
Figure 12: Recovery under Repeated Perturbations: (b)-(c) externally induced second fall; (d)–(g) unstable stand-up attempt leading to a re-fall; (h)–(j) autonomous second recovery.The Video is provided on the website.
IV-F Recovery Adaptability on Unseen Terrains
@@
    Random Boxes: Discontinuous platforms with 0.05m0.05m-$0.2m$ height variations.
@@
    Rough Terrain: High-frequency unevenness ($\delta=0.02m$-$0.10m$).
@@
    Sloped Pyramid: Inclined surfaces with $20\%$$60\%$ gradients.
@@
    Pyramid Stairs: Ascending/descending stairs with step heights 0.05m0.05m0.23m0.23m.
@@
Notably, the policy exhibits partial zero-shot generalization: the robot maintains stability through controlled wheel-leg coordination, despite the absence of explicit terrain sensing. We compared the recovery success rates of the ED strategy and the baseline strategy deployed on the KYON robot in non-flat environments, defining recovery as having a joint angle deviation from the default configuration of no more than $0.5rad$ and a base orientation error—calculated as the sum of differences between the robot’s projected gravity vector and the ideal direction $\left[0,0,-1\right]$ —of less than $0.1$, with the success rates shown in the Table IV.
However, recovery success rates decline significantly under these conditions, with frequent secondary falls observed—particularly on stair terrains. These undesirable behaviors, which undermine real-world deployment, can be attributed to two unresolved challenges. First, the policy’s reactive adjustments lack explicit perception of terrain geometry, causing delayed or inadequate responses when contact conditions change abruptly such edge or step transitions. Second, the learned posture adaptation remains largely configuration-driven rather than terrain-aware, resulting in suboptimal body alignment and insufficient stability margins on uneven or sloped surfaces.
@@
Figure 13: Evaluation of the ED-policy on the KYON Robot in irregular terrain.
@@ V CONCLUSION
This paper presents a Episode-based Dynamic Reward Shaping for wheel-legged robots to achieve robust and adaptive recovery from fallen states. By combining dynamic reward shaping and curriculum learning, the proposed method effectively overcomes the exploration-convergence dilemma inherent in sparse-reward reinforcement learning. Experimental results demonstrate that ED achieves a recovery success rate of over $97\%$ across diverse platforms, outperforming baseline methods by 3 percentage points. The core innovation stems in the synergy between wheel-assisted motion and leg articulation: wheels facilitate rapid centroidal adjustments via controlled rolling, while legs ensure fine-grained posture stabilization., collectively reducing joint torque demands by $15-26\%$ compared to fixed-wheel configurations. Cross-platform validation on robots with distinct kinematic and dynamic parameters (KYON and Unitree Go2-W[8]) further confirms the generalization capability of the framework, where ED autonomously adapts to hardware-specific constraints without manual tuning. These findings highlight the potential of wheel-leg coordination as a universal strategy for recovery control in hybrid locomotion systems.
Future work will focus on extending the framework to handle more complex terrains (e.g., slopes, gravel) and integrating real-world sensor noise models to bridge the sim-to-real gap. Moreover, the principles of dynamic reward shaping could be generalized to other discontinuous-contact tasks, including multi-robot collaborative recovery, manipulation-oriented posture adjustment, and adaptive locomotion in cluttered environments, potentially enabling more robust, versatile, and energy-efficient robotic behaviors.
@@ ACKNOWLEDGMENT
We would like to thank Andrea Patrizi, Despoina Maligianni, Rui Dai, Yifang Zhang, Maolin Lei, Jingcheng Jiang, Kuanqi Cai, and Carlo Rizzardo for their useful discussions.
References Section
  • [1] C. D. Bellicoso, M. Bjelonic, L. Wellhausen, K. Holtmann, F. Günther, M. Tranzatto, P. Fankhauser, and M. Hutter, “Advances in real-world applications for legged robots,” Journal of Field Robotics, vol. 35, no. 8, pp. 1311–1326, 2018.
  • [2] M. Bjelonic, V. Klemm, J. Lee, and M. Hutter, “A survey of wheeled-legged robots,” in Climbing and Walking Robots Conference. Springer, 2022, pp. 83–94.
  • [3] A. Bouman, M. F. Ginting, N. Alatur, M. Palieri, D. D. Fan, T. Touma, T. Pailevanian, S.-K. Kim, K. Otsu, J. Burdick et al., “Autonomous spot: Long-range autonomous exploration of extreme environments with legged locomotion,” in 2020 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS). IEEE, 2020, pp. 2518–2525.
  • [4] M. Tranzatto, F. Mascarich, L. Bernreiter, C. Godinho, M. Camurri, S. Khattak, T. Dang, V. Reijgwart, J. Loeje, D. Wisth et al., “Cerberus: Autonomous legged and aerial robotic exploration in the tunnel and urban circuits of the darpa subterranean challenge,” arXiv preprint arXiv:2201.07067, vol. 3, 2022.
  • [5] H. Christensen, N. Amato, H. Yanco, M. Mataric, H. Choset, A. Drobnis, K. Goldberg, J. Grizzle, G. Hager, J. Hollerbach et al., “A roadmap for us robotics–from internet to robotics 2020 edition,” Foundations and Trends® in Robotics, vol. 8, no. 4, pp. 307–424, 2021.
  • [6] Y. Ma, F. Farshidian, and M. Hutter, “Learning arm-assisted fall damage reduction and recovery for legged mobile manipulators,” in 2023 IEEE International Conference on Robotics and Automation (ICRA). IEEE, 2023, pp. 12 149–12 155.
  • [7] J. A. Castano, C. Zhou, and N. Tsagarakis, “Design a fall recovery strategy for a wheel-legged quadruped robot using stability feature space,” in 2019 IEEE International Conference on Robotics and Biomimetics (ROBIO). IEEE, 2019, pp. 41–46.
  • [8] U. Robotics, “Unitree go2-w,” 2024. [Online]. Available: https://github.com/unitreerobotics/unitree_ros/tree/master/robots/go2w_description
  • [9] C. Semini, J. Goldsmith, B. U. Rehman, M. Frigerio, V. Barasuol, M. Focchi, and D. G. Caldwell, “Design overview of the hydraulic quadruped robots,” in The fourteenth Scandinavian international conference on fluid power. sn, 2015, pp. 20–22.
  • [10] U. Saranli, A. A. Rizzi, and D. E. Koditschek, “Model-based dynamic self-righting maneuvers for a hexapedal robot,” The International Journal of Robotics Research, vol. 23, no. 9, pp. 903–918, 2004.
  • [11] J. Lee, J. Hwangbo, and M. Hutter, “Robust recovery controller for a quadrupedal robot using deep reinforcement learning,” arXiv preprint arXiv:1901.07517, 2019.
  • [12] J. Hwangbo, J. Lee, A. Dosovitskiy, D. Bellicoso, V. Tsounis, V. Koltun, and M. Hutter, “Learning agile and dynamic motor skills for legged robots,” Science Robotics, vol. 4, no. 26, p. eaau5872, 2019.
  • [13] M. Mittal, C. Yu, Q. Yu, J. Liu, N. Rudin, D. Hoeller, J. L. Yuan, R. Singh, Y. Guo, H. Mazhar, A. Mandlekar, B. Babich, G. State, M. Hutter, and A. Garg, “Orbit: A unified simulation framework for interactive robot learning environments,” IEEE Robotics and Automation Letters, vol. 8, no. 6, pp. 3740–3747, 2023.
  • [14] T. Xiao, E. Jang, D. Kalashnikov, S. Levine, J. Ibarz, K. Hausman, and A. Herzog, “Thinking while moving: Deep reinforcement learning with concurrent control,” arXiv preprint arXiv:2004.06089, 2020.
  • [15] C. Zhang, W. Yu, and Z. Li, “Accessibility-based clustering for efficient learning of locomotion skills,” in 2022 International Conference on Robotics and Automation (ICRA). IEEE, 2022, pp. 1600–1606.
  • [16] J. Lee, J. Hwangbo, L. Wellhausen, V. Koltun, and M. Hutter, “Learning quadrupedal locomotion over challenging terrain,” Science robotics, vol. 5, no. 47, p. eabc5986, 2020.
  • [17] N. Rudin, D. Hoeller, P. Reist, and M. Hutter, “Learning to walk in minutes using massively parallel deep reinforcement learning,” in Conference on Robot Learning. PMLR, 2022, pp. 91–100.